OpenStudy (anonymous):

Please Help!!!!!!!! Solve the equations for the solutions, x and y . x + y = k 2x + 3y = k + 1 The lines intersect at __, __

OpenStudy (eamier):

you need to solve the equation for x and y in term of k

OpenStudy (eamier):

make the coefficient of x in eq.1 same as eq2 then eq2-eq1. can you understand what i am saying

OpenStudy (eamier):

you will get the y value

OpenStudy (wolf1728):

A) x + y = k B) 2x + 3y = k + 1 Multiply equation A by -2 A) -2x -2y =-2k now add this to B B) 2x + 3y = k + 1 y = -k + 1 Multiply equation A by -3 A) -3x -3y = -3k now add this to B B) 2x + 3y = k + 1 -x = -2k + 1 or x = 2k -1